# Exploring the magnetic landscape of easily-exfoliable two-dimensional materials

### v1

@@ Abstract
Magnetism in two-dimensional (2D) van der Waals (vdW) materials has garnered significant attention in recent years, 1; 2; 3; 4; 5; 6; 7; 8; 9 not only for potential technological applications but also for the large variety of phenomena that can be hosted by magnetic 2D materials. Indeed, magnetism in 2D displays very different properties depending on the interplay between thermal fluctuations, exchange interactions, and magnetic anisotropy. The thrust to investigate this richness in experiments has stimulated an intense search for novel magnetic monolayers displaying specific phenomena, from Ising anisotropy, 10 to Berezinskii-Kosterlitz-Thouless transitions, 11 and heavy-fermion excitations 12.
@@
Although many candidates have been suggested by a long-term experience with layered magnetic materials, the continuous demand for novel 2D magnets calls for more systematic and exhaustive investigations. In this respect, theoretical approaches, particularly first-principles calculations based on density functional theory (DFT), provide a robust and high-throughput approach to predict novel 2D materials, 13; 14; 15; 16; 17; 18; 19; 20; 21; 22 including magnetic monolayers 13; 23; 24 and their properties. 25; 26
@@
Despite many successes, DFT can sometimes struggle to accurately predict the magnetic properties of materials, owing to the localized nature of - and -orbitals typically involved in magnetism and their tendency to suffer from self-interaction errors when adopting approximate exchange-correlation functionals. A computationally inexpensive and effective strategy to overcome these limitations is achieved by complementing approximate functionals with self-interaction corrections 27 in the so-called Hubbard DFT+ approach. 28; 29; 30 The Hubbard parameter entering the DFT+ approach can be considered, somehow unsatisfactory, as a semi-empirical parameter, or can be computed self-consistently, e.g. using linear-response theory 31. In the latter case, the theory preserves a first-principles character, and efficient implementations using density-functional perturbations theory (DFPT) are suitable for high-throughput investigations. 32; 33; 34; 35; 31; 36; 30; 37; 38
@@
Another crucial challenge for simulations is the identification of the ground state magnetic configuration of the system. Indeed, the magnetic energy landscape is typically complex, with several local minima associated with different magnetic states. The inclusion of Hubbard corrections, needed for a better description of magnetic systems, typically leads to a proliferation of local minima, making it even more challenging to identify the true ground state (global minimum).39; 40; 41; 42; 43; 44; 45; 46; 47; 48 Several approaches have been developed to tackle this issue, ranging from efficiently generating different magnetic configurations, 49 to the manipulation of the occupation matrix for - or -orbitals, 50; 51; 52; 53; 39 to computing spin waves 54 and spin spirals 55, employing cluster multiple theory, 56 or even genetic algorithms. 57
@@
In this work, we thus explore the magnetic properties of easily exfoliable 2D materials from the Materials Cloud two-dimensional crystals database (MC2D) 14; 13; 58; 59 using our recently developed approaches to reveal novel magnetic monolayers. Starting from 3077 easily and potentially exfoliabble monolayers (2004 easily and 1073 potentially exfoliable) with up to 40 atoms per cell 14; 13, 877 systems with up to 12 atoms per cell are screened with plain DFT using an AiiDA60; 61 workflow (Chronos)13. Out of these systems, 483 monolayers are found to be non-magnetic, 166 systems were discarded because of the failure in the optimization, and 228 monolayers showed a magnetic ground state. For these 228 materials, Hubbard corrections are then included for better accuracy, with the Hubbard parameter computed self-consistently. To identify the magnetic ground state, we employ the Robust Occupation Matrix Energy Optimization (RomeoDFT) algorithm 48 based on constraining the atomic orbital occupation matrices that allows us to systematically explore the magnetic energy landscape. With this protocol, we identify 194 magnetic 2D materials for which we accurately determine the magnetic ground state and investigate electronic and magnetic properties, including a proxy for the critical temperature. Notably, we identify 12 half-metals, a class of materials with spin-dependent conductivity, where electrons in one spin channel (up or down) are metallic, while those in the opposite channel are insulating, which is especially promising for spintronics applications. The remaining 34 systems were discarded because either they were repeated, found non-magnetic, consistently failed in the calculation of the Hubbard parameter, or not enough minima were found using RomeoDFT (see Supporting Information).
@@
To identify magnetic 2D materials and their magnetic ground state, we consider monolayers from MC2D,58; 59 a portfolio of 2D materials that can be exfoliated from experimentally known 3D parent compounds, identified using high-throughput computational exfoliation. 13; 14 We start from 877 easily exfoliable 2D materials with up to 12 atoms per unit cell. As summarized in Fig. 1, the tendency to show magnetism is first tested using an improved version of the AiiDA 60; 61 workflow, named Chronos, originally adopted in Ref. 13 to screen a smaller set of materials. For each system, the workflow first considers several ferromagnetic configurations through different starting magnetizations on the atoms and computes its optimized geometry and total energy through collinear DFT calculations with an ordinary approximate exchange-correlation functional (PBE). If any ferromagnetic state is found to have a total energy lower than a non-magnetic reference calculation, the system is considered potentially magnetic and additional magnetic configurations are investigated, including antiferromagnetic states that require up to a supercell to accommodate two magnetic atoms with opposite spins (see Methods for more details). In this way, 228 magnetic monolayers are identified with up to 12 atoms in their primitive cell, together with their potential collinear ground state.
@@
Although versatile and flexible, the workflow explores possible magnetic states only through the starting magnetization, that is, through initial conditions on the density of spin up and spin down states on each atom. A more robust way is to initialize from the atomic orbital occupation matrices - particularly and orbitals where the magnetization usually stems from - and explore more possible states. Introducing atomic orbitals labeled by on the -th atom, we can define an atomic occupation matrix as , where are the occupations of Bloch states and magnetization is an unbalance between the occupation of spin-up and down states . To explore more systematically the magnetic energy landscape and obtain a more reliable identification of the magnetic ground state, it is important to gain full control over the atomic occupation matrix (at least for the relevant atomic orbitals on the magnetic atoms), not just through the spin density. 50; 51; 52; 53; 39 We thus validate the magnetic ground-state identification by applying a constraint on the atomic orbital occupations and then letting the self-consistent procedure find the closest local minimum in the energy landscape. The overall global minimum corresponding to the magnetic ground state can then be found by exhaustively exploring possible atomic occupations and comparing the energies of the corresponding self-consistent solutions. To perform this task automatically and efficiently, we use the RomeoDFT workflow 48 (see Methods for more details). We first consider a subgroup of magnetic monolayers and find that without Hubbard corrections in essentially all cases the same ground state is found either by controlling the starting magnetization by spin density or the constraints on the full atomic occupation matrix (see figure S1 in Supporting Information. When considering PBE calculations, the search implemented in the Chronos workflow is thus exhaustive enough to reproduce the more sophisticated exploration performed by RomeoDFT.
@@
Most of the identified 2D magnetic materials host electrons in localized or orbitals, for which approximate DFT functionals such as PBE tend to suffer from inaccuracies arising from self-interaction. 27; 62 The magnetic ground-state predictions might be affected by such errors and need to be validated against more accurate calculations. As alluded to before, an approach to mitigate self-interaction problems, known as DFT+, relies on complementing DFT functionals with terms inspired by the Hubbard model. 63; 64; 65; 28; 66; 67; 68; 30; 69 Such Hubbard-corrected functionals have been shown to remove prominent self-interaction errors when the Hubbard parameter is computed self-consistently from linear-response theory, 31 thus preserving the parameter-free character of DFT. Unfortunately, the inclusion of Hubbard corrections makes the energy landscape much more complex, with the appearance of many emergent local minima with similar total energies. 48 To illustrate this effect, we consider the case of monolayer CoO2 in Fig. 2, where each point represents a local minimum in the energy landscape, identified by exploring multiple occupation matrices. At the PBE level (), only a limited set of local minima is found, with an overall ferromagnetic ground state (indicated by a star). When including Hubbard corrections, with the Hubbard calculated self-consistently from the linear-response method (PBE+), the number of local minima significantly increases, making the magnetic energy landscape more complex and the possibility of falling in a local minimum different from the true ground state higher. We thus have that, while at the PBE level () the number of states is limited and the identification of the global ground-state minimum is simpler (explaining the agreement between Chronos and RomeoDFT at the PBE level (figure S1 in the Supporting information)), an approach based on the occupation matrix is mandatory when Hubbard corrections are included. In Fig. 2 we also show results for an intermediate case with eV, where the proliferation of local minima is already present, although less dramatic than at . Remarkably, the ground state, labeled by a star in each panel, is very sensitive to the value of as the system is predicted to be ferromagnetic at , antiferromagnetic at 4 eV and then ferromagnetic again (but with a different magnetization) at eV. 70 Therefore, the choice of strongly affects the ground-state properties of the system. To avoid any empiricism, we compute the Hubbard parameter for each material using an efficient formulation of the linear-response approach 31 within density-functional perturbation theory 33; 34; 35 (see Methods and Supporting Information for more details).
@@
Equipped with self-consistent Hubbard corrections, we navigate the magnetic energy landscape of potentially 2D magnetic materials identified through the Chronos workflow at the PBE level, now using PBE+ and a control over the atomic occupation matrix using RomeoDFT. Out of 228 materials, 34 systems are discarded because either they are found to be nonmagnetic or there are convergence issues in the calculation of the Hubbard parameter or self-consistent calculations, leaving us with 194 magnetic monolayers for which we characterize their ground state. The results are summarized in Fig. 3, where we report the absolute magnetization per magnetic atom and the band gap of the ground state of all systems, together with their distribution. More details on the specific magnetic configuration, crystal structure, magnetic energy landscape, and band structure for each material are provided in the Supporting Information.
@@
fig. 4 shows the number of systems in each group, with the left and right charts representing results from PBE (Chronos workflow) and PBE+ (RomeoDFT workflow) calculations, respectively. Blue and red colors indicate antiferromagnetic and ferromagnetic materials. The innermost rings show the number of systems in each category, while the second rings count metallic (hatched) or semiconducting (solid) systems. In the right chart, the outermost ring indicates the number of systems from the innermost ring that either remained in the same magnetic category (solid fill) or changed categories (dotted fill) from the PBE calculations. As expected, the inclusion of Hubbard corrections leads to an increase in the number of insulating systems. From fig. 4, we also note that using the RomeoDFT workflow with PBE+ calculations significantly expands the number of antiferromagnetic systems (blue colors): approximately 40% of these materials were identified as ferromagnetic when using the Chronos workflow with standard PBE calculations. This change can be attributed primarily to two factors: first, the limited capability to explore the energy landscape when using simply a starting magnetization; second, the possible increase in the number of magnetic states after including Hubbard corrections.
@@
We note in passing that among the AFM identified above, we find only two altermagnetic monolayers by generalizing the protocol for 3D systems 71 to include additional symmetry constraints in 2D. 72; 73 These are RuF4 and VF4, with the same structural prototype. Both systems were already predicted to be 2D altermagnets in Ref. 72; we do not see altermagnetism in FeBr3, because its crystal structure and symmetry is different from the one studied in Ref. 72.
@@
Due to the scarcity of experimental results for magnetic monolayers, it is challenging to validate our predictions against experiments. We thus focus on a few cases reported in the existing literature. For example, our results agree with the experimentally observed ferromagnetism in the family of chromium trihalides (CrX3; X = Cl, Br, I with CrCl3 being 2D-XY ferromagnetic) 74; 10; 11 and in CrSBr. 75
@@
In the family of transition-metal phosphorus trisulfides, our results match the experimentally observed ferromagnetism in CrPS4 76 and the antiferromagnetic ordering in NiPS3, although suppressed by fluctuations in the monolayer limit. 77 For the case of FePS3, our findings show ferromagnetic ordering, in contrast with the antiferromagnetic ordering observed in experiments. 78 We attribute this discrepancy to the limited size of the unit cell considered in this work for FePS3, which is not able to accommodate the measured zigzag ordering. Indeed, since the primitive cell already contains two magnetic atoms, no supercell is considered, and only a Néel AFM ordering is tested. This limitation, combined with the use of collinear calculations, could influence the accuracy in predicting other antiferromagnetic systems and may lead to discrepancies with other studies, 79; 55, and we expect it to be at the origin of the relatively small number of AFM systems with respect to FM.
@@
Among the possible conclusions that we can draw from the high-throughput investigation of the magnetic ground-state of 2D materials, we also focus on the nature of magnetic elements in each structure. Conventional magnetism usually arises from or orbitals, qualifying transition metals and lanthanides as the dominant contributors to magnetism. Our results largely confirm such conventional wisdom, as shown in fig. 5, where we report the number of systems in which a given element is magnetic. However, fig. 5 also suggests that certain halogen and nonmetal elements (especially oxygen) also play a role in magnetism in 2D compounds, putting forward their potential role in unconventional magnetism. For instance, in GaSe (space group: ) and SiN2F6, the selenium and nitrogen atoms exhibit non-negligible magnetic moments (0.2 and 2 respectively). Interestingly, in the case of CdOCl, while the orbitals of Cd are fully occupied and exhibit negligible magnetic moments, magnetism predominantly arises from the oxygen atoms, which have a magnetic moment of 0.8 .
@@
fig. 6 shows divided by the number of magnetic atoms per cell, , for the insulating magnetic monolayers for which it was possible to find a metastable magnetic state with similar atomic magnetic moments but different ordering. Systems are sorted with increasing effective . The bottom panel contains systems where exchange interactions are weak, making it more challenging for the magnetic ordering to compete with thermal fluctuations. On the contrary, the systems in the top panel are expected to have higher critical transition temperatures. The ranking based on the effective seems to reproduce several experimental trends for the ordering temperature. For instance, it is experimentally known that CrSBr has a higher critical temperature than the chromium trihalides, 10; 75 and within this family the iodine compound has the largest while the chlorine one the smallest, 80; 81; 82; 10; 11 in agreement with the ordering in fig. 6.
@@
To further validate as a meaningful proxy for , we select three systems with the largest in fig. 6 and compute their transition temperatures using different approaches. We thus focus on MoCl3, GdI2, and VPS3, noting that the latter was already identified as a potentially high- magnetic monolayer based on high-throughput calculations. 83 Additionally, we include CrI3 in our analysis, as experimental results are available for comparison. In addition to the approach based on eq. 1 and relying on the energy difference between FM and AFM reference states, we also consider the so-called four-state method, 84 which calculates the exchange parameter based on the energy differences between four distinct magnetic configurations of a pair of neighboring atoms in a supercell (see Methods for more details). Remarkably, the four-state method allows the calculation of specific interatomic exchange parameters, not just an effective nearest-neighbor as in eq. 1, and we consider interactions up to the third (fourth) nearest neighbors for CrI3 and GdI2 (VPS3 and MoCl3). The exchange parameters computed within the two approaches are then denoted and in the following. From the knowledge of the exchange parameters, the transition temperature is calculated using either mean-field (MF) theory or classical Monte Carlo (MC) simulations. In the latter, we completely neglect relativistic effects. Indeed, magnetocrystalline anisotropy does not affect significantly the critical temperature for a finite-size system, and even isotropic monolayers can sustain magnetic order at finite temperature when considering realistic sample sizes. 85
@@
Critical temperatures computed either using MF or MC, with exchange parameters estimated from both eq. 1 and the four-state method, are shown in fig. 7 for CrI3, GdI2, VPS3, and MoCl3. Trends in derived from both MC and MF methods are consistent with fig. 6, with MF calculations showing higher values, as expected from the well-known overestimation of in mean-field theory. 86
@@
We note that, in general, different methods for computing the exchange parameters give similar values of . The only exception is MoCl3, for which the MC estimates of differ significantly between the two methods, with a notably lower value when the exchange parameters are calculated using the four-state method. This discrepancy is attributed to the dimerization of Mo atoms in MoCl3, which distorts the otherwise honeycomb lattice of Mo atoms. The energy difference between the FM and AFM is mainly arising from a strong AFM coupling between dimerized pairs, while dimers are only weakly coupled. This feature is correctly captured within the four-state method, while the effective approach of eq. 1 uniformly distributes the FM-AFM energy difference between the closest atoms. This failure also affects all MF estimates in a similar way. We thus expect that the effective reported in fig. 6 provides a good estimation of the ranking in critical temperature, provided that exchange interactions are distributed evenly between similar neighbors.
@@
We also note that the transition temperature of the CrI3 monolayer is higher than the values reported in previous studies and experiments87; 25; 10. This discrepancy can be attributed to differences in geometry, the relatively large value for CrI3 obtained from linear-response,38, and the influence of broadening effects in Brillouin-zone integrations. For consistency, we employed the same broadening value as used in the Chronos workflow (270 meV). Since the exchange parameter is derived from total energies–which are sensitive to the choice of broadening–it is likewise influenced by the selected broadening value.

### v2

@@ Abstract
two-dimensional magnetic materials, hubbard corrections, global and local minima, energy landscape, density-functional theory, DFT+Magnetism in two-dimensional (2D) van der Waals (vdW) materials has garnered significant attention in recent years, 1, 2, 3, 4, 5, 6, 7, 8, 9 not only for potential technological applications but also for the large variety of phenomena that can be hosted by magnetic 2D materials. Indeed, magnetism in 2D displays very different properties depending on the interplay between thermal fluctuations, exchange interactions, and magnetic anisotropy. The thrust to investigate this richness in experiments has stimulated an intense search for novel magnetic monolayers displaying specific phenomena, from Ising anisotropy, 10 to Berezinskii-Kosterlitz-Thouless transitions, 11 and heavy-fermion excitations 12.
@@
Although many candidates have been suggested by a long-term experience with layered magnetic materials, the continuous demand for novel 2D magnets calls for more systematic and exhaustive investigations. In this respect, theoretical approaches, particularly first-principles calculations based on density functional theory (DFT), provide a robust and high-throughput approach to predict novel 2D materials, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22 including magnetic monolayers 13, 23, 24 and their properties. 25, 26
@@
Despite many successes, DFT can sometimes struggle to accurately predict the magnetic properties of materials, owing to the localized nature of - and -orbitals typically involved in magnetism and their tendency to suffer from self-interaction errors when adopting approximate exchange-correlation functionals. A computationally inexpensive and effective strategy to overcome these limitations is achieved by complementing approximate functionals with self-interaction corrections 27 in the so-called Hubbard DFT+ approach. 28, 29, 30 The Hubbard parameter entering the DFT+ approach can be considered, somehow unsatisfactory, as a semi-empirical parameter, or can be computed self-consistently, e.g. using linear-response theory 31. In the latter case, the theory preserves a first-principles character, and efficient implementations using density-functional perturbations theory (DFPT) are suitable for high-throughput investigations. 32, 33, 34, 35, 31, 36, 30, 37, 38
@@
Another crucial challenge for simulations is the identification of the ground state magnetic configuration of the system. Indeed, the magnetic energy landscape is typically complex, with several local minima associated with different magnetic states. The inclusion of Hubbard corrections, needed for a better description of magnetic systems, typically leads to a proliferation of local minima, making it even more challenging to identify the true ground state (global minimum).39, 40, 41, 42, 43, 44, 45, 46, 47, 48 Several approaches have been developed to tackle this issue, ranging from efficiently generating different magnetic configurations, 49 to the manipulation of the occupation matrix for - or -orbitals, 50, 51, 52, 53, 39 to computing spin waves 54 and spin spirals 55, employing cluster multiple theory, 56 genetic algorithms 57 or machine learning methods.58
@@
In this work, we thus explore the magnetic properties of easily exfoliable 2D materials from the Materials Cloud two-dimensional crystals database (MC2D) 14, 13, 59, 60 using our recently developed approaches to reveal novel magnetic monolayers. Starting from 3077 easily and potentially exfoliabble monolayers (2004 easily and 1073 potentially exfoliable) with up to 40 atoms per cell 14, 13, 877 systems with up to 12 atoms per cell are screened with plain DFT using an AiiDA61, 62 workflow (Chronos)13. Out of these systems, 483 monolayers are found to be non-magnetic, 166 systems were discarded because of the failure in the optimization, and 228 monolayers showed a magnetic ground state. For these 228 materials, Hubbard corrections are then included for better accuracy, with the Hubbard parameter computed self-consistently. To identify the magnetic ground state, we employ the Robust Occupation Matrix Energy Optimization (RomeoDFT) algorithm 48 based on constraining the atomic orbital occupation matrices that allows us to systematically explore the magnetic energy landscape. With this protocol, we identify 194 magnetic 2D materials for which we accurately determine the magnetic ground state and investigate electronic and magnetic properties, including a proxy for the critical temperature. Notably, we identify 12 half-metals, a class of materials with spin-dependent conductivity, where electrons in one spin channel (up or down) are metallic, while those in the opposite channel are insulating, which is especially promising for spintronics applications. The remaining 34 systems were discarded because either they were repeated, found non-magnetic, consistently failed in the calculation of the Hubbard parameter, or not enough minima were found using RomeoDFT (see Supporting Information).
@@
To identify magnetic 2D materials and their magnetic ground state, we consider monolayers from MC2D,59, 60 a portfolio of 2D materials that can be exfoliated from experimentally known 3D parent compounds, identified using high-throughput computational exfoliation. 13, 14 We start from 877 easily exfoliable 2D materials with up to 12 atoms per unit cell. As summarized in Figure 1, the tendency to show magnetism is first tested using an improved version of the AiiDA 61, 62 workflow, named Chronos, originally adopted in Ref. 13 to screen a smaller set of materials. For each system, the workflow first considers several ferromagnetic configurations through different starting magnetizations on the atoms and computes its optimized geometry and total energy through collinear DFT calculations with an ordinary approximate exchange-correlation functional (PBE). If any ferromagnetic state is found to have a total energy lower than a non-magnetic reference calculation, the system is considered potentially magnetic and additional magnetic configurations are investigated, including antiferromagnetic states that require up to a supercell to accommodate two magnetic atoms with opposite spins (see Methods for more details). In this way, 228 magnetic monolayers are identified with up to 12 atoms in their primitive cell, together with their potential collinear ground state.
@@
Although versatile and flexible, the workflow explores possible magnetic states only through the starting magnetization, that is, through initial conditions on the density of spin up and spin down states on each atom. A more robust way is to initialize from the atomic orbital occupation matrices - particularly and orbitals where the magnetization usually stems from - and explore more possible states. Introducing atomic orbitals labeled by on the -th atom, we can define an atomic occupation matrix as , where are the occupations of Bloch states and magnetization is an unbalance between the occupation of spin-up and down states . To explore more systematically the magnetic energy landscape and obtain a more reliable identification of the magnetic ground state, it is important to gain full control over the atomic occupation matrix (at least for the relevant atomic orbitals on the magnetic atoms), not just through the spin density. 50, 51, 52, 53, 39 We thus validate the magnetic ground-state identification by applying a constraint on the atomic orbital occupations and then letting the self-consistent procedure find the closest local minimum in the energy landscape. The overall global minimum corresponding to the magnetic ground state can then be found by exhaustively exploring possible atomic occupations and comparing the energies of the corresponding self-consistent solutions. To perform this task automatically and efficiently, we use the RomeoDFT workflow 48 (see Methods for more details). We first consider a subgroup of magnetic monolayers and find that without Hubbard corrections in essentially all cases the same ground state is found either by controlling the starting magnetization by spin density or the constraints on the full atomic occupation matrix (see figure S1 in Supporting Information. When considering PBE calculations, the search implemented in the Chronos workflow is thus exhaustive enough to reproduce the more sophisticated exploration performed by RomeoDFT.
@@
Most of the identified 2D magnetic materials host electrons in localized or orbitals, for which approximate DFT functionals such as PBE tend to suffer from inaccuracies arising from self-interaction. 27, 63 The magnetic ground-state predictions might be affected by such errors and need to be validated against more accurate calculations. As alluded to before, an approach to mitigate self-interaction problems, known as DFT+, relies on complementing DFT functionals with terms inspired by the Hubbard model. 64, 65, 66, 28, 67, 68, 69, 30, 70 Such Hubbard-corrected functionals have been shown to remove prominent self-interaction errors when the Hubbard parameter is computed self-consistently from linear-response theory, 31 thus preserving the parameter-free character of DFT. Unfortunately, the inclusion of Hubbard corrections makes the energy landscape much more complex, with the appearance of many emergent local minima with similar total energies. 48 To illustrate this effect, we consider the case of monolayer CoO2 in Figure 2, where each point represents a local minimum in the energy landscape, identified by exploring multiple occupation matrices. At the PBE level (), only a limited set of local minima is found, with an overall ferromagnetic ground state (indicated by a star). When including Hubbard corrections, with the Hubbard calculated self-consistently from the linear-response method (PBE+), the number of local minima significantly increases, making the magnetic energy landscape more complex and the possibility of falling in a local minimum different from the true ground state higher. We thus have that, while at the PBE level () the number of states is limited and the identification of the global ground-state minimum is simpler (explaining the agreement between Chronos and RomeoDFT at the PBE level — see fig. S1 in the Supporting information), an approach based on the occupation matrix is mandatory when Hubbard corrections are included. In Figure 2 we also show results for an intermediate case with eV, where the proliferation of local minima is already present, although less dramatic than at . Remarkably, the ground state, labeled by a star in each panel, is very sensitive to the value of as the system is predicted to be ferromagnetic at , antiferromagnetic at 4 eV and then ferromagnetic again (but with a different magnetization) at eV. 71 Therefore, the choice of strongly affects the ground-state properties of the system. To avoid any empiricism, we compute the Hubbard parameter for each material using an efficient formulation of the linear-response approach 31 within density-functional perturbation theory 33, 34, 35 (see Methods and Supporting Information for more details).
@@
Equipped with self-consistent Hubbard corrections, we navigate the magnetic energy landscape of potentially 2D magnetic materials identified through the Chronos workflow at the PBE level, now using PBE+ and a control over the atomic occupation matrix using RomeoDFT. Out of 228 materials, 34 systems are discarded because either they are found to be nonmagnetic or there are convergence issues in the calculation of the Hubbard parameter or self-consistent calculations, leaving us with 194 magnetic monolayers for which we characterize their ground state. The results are summarized in Figure 3, where we report the absolute magnetization per magnetic atom and the band gap of the ground state of all systems, together with their distribution. More details on the specific magnetic configuration, crystal structure, magnetic energy landscape, and band structure for each material are provided in the Supporting Information.
@@
Figure 4 shows the number of systems in each group, with the left and right charts representing results from PBE (Chronos workflow) and PBE+ (RomeoDFT workflow) calculations, respectively. Blue and red colors indicate antiferromagnetic and ferromagnetic materials. The innermost rings show the number of systems in each category, while the second rings count metallic (hatched) or semiconducting (solid) systems. In the right chart, the outermost ring indicates the number of systems from the innermost ring that either remained in the same magnetic category (solid fill) or changed categories (dotted fill) from the PBE calculations. As expected, the inclusion of Hubbard corrections leads to an increase in the number of insulating systems. From Figure 4, we also note that using the RomeoDFT workflow with PBE+ calculations significantly expands the number of antiferromagnetic systems (blue colors): approximately 40% of these materials were identified as ferromagnetic when using the Chronos workflow with standard PBE calculations. This change can be attributed primarily to two factors: first, the limited capability to explore the energy landscape when using simply a starting magnetization; second, the possible increase in the number of magnetic states after including Hubbard corrections.
@@
We note in passing that among the AFM identified above, we find only two altermagnetic monolayers by generalizing the protocol for 3D systems 72 to include additional symmetry constraints in 2D. 73, 74 These are RuF4 and VF4, with the same structural prototype. Both systems were already predicted to be 2D altermagnets in Ref. 73; we do not see altermagnetism in FeBr3, because its crystal structure and symmetry is different from the one studied in Ref. 73.
@@
Due to the scarcity of experimental results for magnetic monolayers, it is challenging to validate our predictions against experiments. We thus focus on a few cases reported in the existing literature. For example, our results agree with the experimentally observed ferromagnetism in the family of chromium trihalides (CrX3; X = Cl, Br, I with CrCl3 being 2D-XY ferromagnetic) 75, 10, 11 and in CrSBr. 76
@@
In the family of transition-metal phosphorus trisulfides, our results match the experimentally observed ferromagnetism in CrPS4 77 and the antiferromagnetic ordering in NiPS3, although suppressed by fluctuations in the monolayer limit. 78 For the case of FePS3, our findings show ferromagnetic ordering, in contrast with the antiferromagnetic ordering observed in experiments. 79 We attribute this discrepancy to the limited size of the unit cell considered in this work for FePS3, which is not able to accommodate the measured zigzag ordering. Indeed, since the primitive cell already contains two magnetic atoms, no supercell is considered, and only a Néel AFM ordering is tested. This limitation, combined with the use of collinear calculations, could influence the accuracy in predicting other antiferromagnetic systems and may lead to discrepancies with other studies, 80, 55 and we expect it to be at the origin of the relatively small number of AFM systems with respect to FM.
@@
Among the possible conclusions that we can draw from the high-throughput investigation of the magnetic ground-state of 2D materials, we also focus on the nature of magnetic elements in each structure. Conventional magnetism usually arises from or orbitals, qualifying transition metals and lanthanides as the dominant contributors to magnetism. Our results largely confirm such conventional wisdom, as shown in Figure 5, where we report the number of systems in which a given element is magnetic. However, Figure 5 also suggests that certain chalcogen and nonmetal elements (especially oxygen) also play a role in magnetism in 2D compounds, putting forward their potential role in unconventional magnetism. For instance, in GaSe (space group: ) and SiN2F6, the selenium and nitrogen atoms exhibit non-negligible magnetic moments (0.2 and 2 respectively). Interestingly, in the case of CdOCl, while the orbitals of Cd are fully occupied and exhibit negligible magnetic moments, magnetism predominantly arises from the oxygen atoms, which have a magnetic moment of 0.8 .
@@
Figure 6 shows divided by the number of magnetic atoms per cell, , for the insulating magnetic monolayers for which it was possible to find a metastable magnetic state with similar atomic magnetic moments but different ordering. Systems are sorted with increasing effective . The bottom panel contains systems where exchange interactions are weak, making it more challenging for the magnetic ordering to compete with thermal fluctuations. On the contrary, the systems in the top panel are expected to have higher critical transition temperatures. The ranking based on the effective seems to reproduce several experimental trends for the ordering temperature. For instance, it is experimentally known that CrSBr has a higher critical temperature than the chromium trihalides, 10, 76 and within this family the iodine compound has the largest while the chlorine one the smallest, 81, 82, 83, 10, 11 in agreement with the ordering in Figure 6.
@@
To further validate as a meaningful proxy for , we select three systems with the largest in Figure 6 and compute their transition temperatures using different approaches. We thus focus on MoCl3, GdI2, and VPS3, noting that the latter was already identified as a potentially high- magnetic monolayer based on high-throughput calculations. 84 Additionally, we include CrI3 in our analysis, as experimental results are available for comparison. In addition to the approach based on eq. 1 and relying on the energy difference between FM and AFM reference states, we also consider the so-called four-state method, 85 which calculates the exchange parameter based on the energy differences between four distinct magnetic configurations of a pair of neighboring atoms in a supercell (see Methods for more details). Remarkably, the four-state method allows the calculation of specific interatomic exchange parameters, not just an effective nearest-neighbor as in eq. 1, and we consider interactions up to the third (fourth) nearest neighbors for CrI3 and GdI2 (VPS3 and MoCl3). The exchange parameters computed within the two approaches are then denoted and in the following. From the knowledge of the exchange parameters, the transition temperature is calculated using either mean-field (MF) theory or classical Monte Carlo (MC) simulations. In the latter, we completely neglect relativistic effects. Indeed, magnetocrystalline anisotropy does not affect significantly the critical temperature for a finite-size system, and even isotropic monolayers can sustain magnetic order at finite temperature when considering realistic sample sizes. 86
@@
Critical temperatures computed either using MF or MC, with exchange parameters estimated from both eq. 1 and the four-state method, are shown in Figure 7 for CrI3, GdI2, VPS3, and MoCl3. Trends in derived from both MC and MF methods are consistent with Figure 6, with MF calculations showing higher values, as expected from the well-known overestimation of in mean-field theory. 87
@@
We note that, in general, different methods for computing the exchange parameters give similar values of . The only exception is MoCl3, for which the MC estimates of differ significantly between the two methods, with a notably lower value when the exchange parameters are calculated using the four-state method. This discrepancy is attributed to the dimerization of Mo atoms in MoCl3, which distorts the otherwise honeycomb lattice of Mo atoms. The energy difference between the FM and AFM is mainly arising from a strong AFM coupling between dimerized pairs, while dimers are only weakly coupled. This feature is correctly captured within the four-state method, while the effective approach of eq. 1 uniformly distributes the FM-AFM energy difference between the closest atoms. This failure also affects all MF estimates in a similar way. We thus expect that the effective reported in Figure 6 provides a good estimation of the ranking in critical temperature, provided that exchange interactions are distributed evenly between similar neighbors.
@@
We also note that the transition temperature of the CrI3 monolayer is higher than the values reported in previous studies and experiments88, 25, 10. This discrepancy can be attributed to differences in geometry, the relatively large value for CrI3 obtained from linear-response,38 and the influence of broadening effects in Brillouin-zone integrations. For consistency, we employed the same broadening value as used in the Chronos workflow (270 meV). Since the exchange parameter is derived from total energies–which are sensitive to the choice of broadening–it is likewise influenced by the selected broadening value.